Patty has 20 coins consisting of nickels and dime. If her nickels were dimes and her dimes were nickels, she would have 70 cents more. How much are her coins worth?
(A) $1.15 (B) $1.20 (C) $1.25 (D) $1.30 (E) $1.35

Answer by richwmiller(17219)

n+d=20
5n+10d+70=5d+10n
5n+10d=a
a=115
d=3
n=17
A) $1.15